# Superluminal constraints from ultra–high–energy neutrino events

### v1

@@ Abstract
The $\sim 100\,$PeV neutrino detected by KM3NeT marks the beginning of ultra–high–energy neutrino astronomy and provides a powerful probe of Lorentz Invariance Violation (LIV). In superluminal scenarios, neutrinos can decay through vacuum $e^{-}e^{+}$ pair emission or neutrino splitting. Previous analyses of the KM3–230213A event relied on simplified survival–probability estimates and, in some cases, used inaccurate decay–width expressions or neglected redshift and threshold effects. In this work we present a unified and self–consistent framework that corrects these issues and applies to both the energy–independent ($n=0$) and quadratic ($n=2$) superluminal cases. We collect and recast the decay–width and threshold expressions, clarify their flavor dependence, and include a consistent treatment of cosmological propagation. We also assess the impact of cascade regeneration and show that cascade effects are negligible for the purpose of setting LIV bounds. The survival–probability approximation adopted in previous works is therefore justified, while our framework provides a coherent basis for future analyses of superluminal neutrino constraints, which should consistently include possible time–delay signatures.
@@ I Introduction
The recent detection of a neutrino event with an energy of approximately 100 PeV by KM3NeT [1] has opened the window of ultra-high-energy (UHE) neutrino astronomy. While its origin—whether it arises from a point source, the diffuse background, or is a cosmogonic neutrino—remains uncertain [2, 3, 4, 5], its mere observation confirms the existence of a population of neutrinos surpassing the 100 PeV threshold [6]. This detection also raises intriguing questions, particularly about the absence of similar events in IceCube, which has so far only reported neutrinos up to an order of magnitude lower in energy [7]. However, it is reasonable to anticipate that this event is just the first of many within this energy regime, to be observed by IceCube, KM3NeT, and next-generation UHE neutrino observatories such as IceCube–Gen2 (radio), RNO–G, GRAND, and POEMMA, which will significantly increase the exposure to neutrinos in the EeV range.
This newly opened window provides an unprecedented opportunity to explore new physics. In fact, the particularities of the KM3-230213A event have been used to connect it with phenomena as diverse as oscillations and sterile neutrinos [8], dark matter [9, 10], and primordial black holes [11, 12].
A population of UHE neutrinos have general implications for new interactions and neutrino decays. In particular, it affects scenarios of Lorentz Invariance Violation (LIV), where neutrinos with a modified dispersion relation may undergo decay during propagation, altering the expected flux of cosmic neutrinos.
@@
The present work revisits these analyses within a unified and self–consistent framework. We incorporate a proper treatment of the decay thresholds, the energy and flavor dependence of the decay widths, and the cosmological propagation of neutrinos. We also examine the validity and limitations of the survival–probability approach by explicitly studying the potential influence of cascade regeneration on the detected flux. This provides a consistent methodology applicable to both energy-independent and energy-dependent superluminal scenarios, and clarifies the domain of applicability of current and future constraints based on ultra-high-energy neutrino observations.
The structure of the paper is as follows. In Sec. II we review the expressions of the decay widths relevant for different LIV scenarios, emphasizing their unified formulation and comparing them with those adopted in the literature. These results are then applied in Sec. III to derive constraints on superluminal neutrinos using the survival–probability method, incorporating the proper redshift dependence and decay thresholds. In Sec. IV we examine the possible impact of cascade regeneration on these results and demonstrate that its contribution is negligible for setting LIV bounds, thereby justifying the validity of the simplified approach. Section V applies this framework to the KM3–230213A event and discusses its implications for future ultra–high–energy neutrino observations, including the connection with possible time–delay effects. Finally, Sec. VI summarizes our conclusions and outlines prospects for future studies.
@@ II Unification of decay widths for superluminal neutrinos
More accurate results for the decay width were obtained in Refs. [19, 20], which showed that the original Cohen–Glashow estimate [17] follows from a prescription that cannot be derived from a consistent Lagrangian framework. More recently, Ref. [18] improved the calculation of both total and partial widths by incorporating the—previously neglected—charged–current contribution to the vacuum $e^{-}e^{+}$ pair–emission (VPE) channel, and by considering that another decay process, $\nu_{\alpha}\to\nu_{\alpha}\nu_{\beta}\bar{\nu}_{\beta}$, neutrino splitting (NSpl), is allowed for $n\neq 0$. The decay widths for the processes $(i)=\{\text{VPE},\text{NSpl}\}$ can be written as
$\Gamma_{\alpha,n}^{(i)}(E)\eqsim K_{\alpha,n}^{(i)}\,\frac{G_{F}^{2}}{192\pi^{3}}\,E^{5+3n}\,\Delta_{n}^{3}\,$, (6)
[…]
@@
$\Gamma_{\alpha,n}^{\text{(All)}}(E)\eqsim\Big(K_{\alpha,n}^{\text{(VPE)}}+K_{\alpha,n}^{\text{(NSpl)}}\Big)\,\frac{G_{F}^{2}}{192\pi^{3}}\,E^{5+3n}\,\Delta_{n}^{3}\,$, (8)
@@
$\frac{m_{e}^{2}/E^{2}}{E^{n}\Delta_{n}}\approx\left(\frac{E_{n}^{\mathrm{th}}}{E}\right)^{n+2}\,.$ (10)
@@
III Constraints on superluminal neutrinos from the survival–probability method
@@
Previous studies constraining superluminal neutrinos [13, 14, 15] have neglected the flavor dependence of the decay widths. In addition, they often rely on the Cohen–Glashow approximation rather than on the full results of Refs. [19, 20, 18], or restrict the analysis to the case $n=0$. In particular, Ref. [13] presented constraints on the LIV parameter $\delta$ for the $n=0$ case, employing the Cohen–Glashow coefficient $1/14\simeq 0.071$ instead of the correct value $17/84\simeq 0.202$ given in Table 1. In addition, the analysis does not incorporate the VPE threshold in their Figure 1, effectively assuming $E>E^{\text{th}}_{n}$, while, as they point out in the main text, the validity of this assumption is limited. Constraints are then derived on $\Delta_{0}=\delta$ as a function of the source distance $L$, using the survival probability $\mathcal{P}^{\text{SV}}_{n}$ introduced in Eq. (11), which strictly applies only to local sources. Nevertheless, this approximation is extended to cosmological distances in their Fig. 1, up to the scale of the observable Universe. In what follows, we address the shortcomings discussed above and develop a framework that overcomes these limitations. Our aim is to provide an analysis scheme that can be applied not only to the KM3-230213A event, but also to future detections of ultra–high–energy neutrinos.
@@
Figure 1: Constraints on the LIV parameter $\delta=\Delta_{0}$ as a function of the source redshift $z$ (bottom axis) and the corresponding source distance $L$ in meters (top axis). The left vertical axis shows the upper bound on $\delta$ obtained from Eqs. (15) and (16) for different detected energies, each represented by a different color as indicated in the legend. The right vertical axis displays the corresponding VPE threshold energy $E^{\text{th}}_{0}$ associated with the values of $\delta$ shown on the left axis.
@@
Figure 2: Heat–map representation of the constraints in the $(E_{d},\delta,z)$ space for the $n=0$ case. The bottom axis shows the LIV parameter, while the top axis gives the corresponding VPE threshold energy, which is also shown as a dashed line in the plot. The left vertical axis corresponds to the detected energy $E_{d}$. Colors encode the constraint derived from the survival–probability condition in the range from $L=10^{17}\,\text{m}$ up to $z=10$. A detected event corresponds to a horizontal line; its intersection with the colored boundary yields the upper bound on $\delta$ at each $z$.
@@
In the $n=0$ case, the curve in Fig. 1 of Ref. [13] is replaced by the set of upper–limit curves we show in Fig. 1. These display the upper bound on $\delta$ as a function of the source redshift $z$ (bottom axis) and the corresponding comoving distance $L$ (top axis). The right vertical axis indicates the VPE threshold energy $E^{\text{th}}_{0}$ corresponding to the value of $\delta$ on the left vertical axis, and each colored curve corresponds to a different detected energy $E_{d}$. Values of $\delta$ above a given curve are excluded for a source at the corresponding $z$ and $E_{d}$. One can also see that the curves are no longer straight lines: the inclusion of cosmic expansion changes their slope at high redshift, and the resulting bound on $\delta$ becomes very weak. In fact, independently of the detected energy $E_{d}$, there always exists a sufficiently distant source ($z$ large enough) for which the survival probability amounts to $\epsilon$. The cosmological effect becomes important already at redshifts $z\sim 1$, which implies that neglecting it, as in Ref. [13], leads to a physically relevant inconsistency in the resulting limits.
@@
A more powerful way of visualizing the constraints is provided by the heat–map shown in Fig. 2, where the three variables $(E_{d},\delta,z)$ are combined into a single representation. Its key advantage is that a detected event corresponds simply to a horizontal line at the measured energy, so that the corresponding upper bound on $\delta$ can be directly read as a function of the source redshift $z$. This representation is therefore valid not only for the present analysis but also for future UHE neutrino detections. The plot spans from a source distance $L=10^{17}\,\text{m}$ (local Universe) up to a redshift $z=10$, and the dashed line indicates the VPE threshold energy associated with each value of $\delta$.
@@
For $n=2$, the relevant parameter to be constrained is not $\delta$, but the LIV scale $\Lambda$, which we express in units of the Planck energy $E_{\text{Pl}}$. Figure 3 shows the lower bound on $\Lambda/E_{\text{Pl}}$ as a function of the source redshift $z$ (bottom axis) and the corresponding source distance $L$ (top axis), for a fixed detected energy $E_{d}$. As in the $n=0$ case, the curves deviate from straight lines at high redshift due to the effect of cosmic expansion, which substantially weakens the constraint.
Figure 3: Constraints on the LIV parameter $\Lambda$ in units of the Planck energy $E_{\text{Pl}}$ as a function of the source redshift $z$ (bottom axis) and the corresponding source distance $L$ (top axis), for fixed detected energy $E_{d}$. The left vertical axis shows the bound on $\Lambda/E_{\text{Pl}}$ obtained from the survival–probability analysis, while the right vertical axis indicates the associated VPE threshold energy. As in the $n=0$ case, the inclusion of cosmic expansion modifies the slope of the curves at large $z$, and the bound on $\Lambda$ rapidly deteriorates.
Figure 4: Heat–map representation of the constraints in the $(E_{d},\Lambda/E_{\text{Pl}},z)$ space for the $n=2$ case. The bottom axis shows the LIV parameter, while the top axis gives the corresponding VPE threshold energy, which is also shown as a dashed line in the plot. The left vertical axis corresponds to the detected energy $E_{d}$. Colors encode the constraint derived from the survival–probability condition in the range from $L=10^{17}\,\text{m}$ up to $z=10$. A detected event corresponds to a horizontal line; its intersection with the colored boundary yields the lower bound on $\Lambda$ at each $z$.
A more event-independent view is provided in Fig. 4, which displays a heat–map representation in the $(E_{d},\Lambda/E_{\text{Pl}},z)$ space. A detected event of energy $E_{d}$ corresponds to a horizontal line, and the intersection with the shaded region directly provides the lower bound on $\Lambda/E_{\text{Pl}}$ as a function of the source redshift. The plot spans the range $L=10^{17}\,\text{m}$ to $z=10$, and the dashed curve marks the corresponding threshold energy.
IV Validity of the survival–probability approximation
The analysis carried out so far has been implicitly based on the assumption that the detected neutrino was directly emitted by the source. In other words, the survival–probability formalism accounts only for the case in which the neutrino has propagated from the source without undergoing any decay. However, in the presence of superluminal propagation, neutrinos may decay along their trajectory, giving rise to a cascade of secondary neutrinos that can also be detected.
@@ IV Validity of the survival–probability approximation
[…]
where $f(y)$ is the distribution of the energy fraction $y$ carried by the secondary neutrino produced in the decay. The first exponential factor accounts for the survival probability of the parent neutrino up to the decay point, the middle factor gives the decay probability within a distance interval dxLdx\,L, and the last exponential factor represents the survival of the daughter neutrino on its way to the detector.
@@
The simplified method introduced in Sec. III, which retains only $\mathcal{R}^{(0)}$, can in principle underestimate the flux of cascade neutrinos. To assess the importance of this effect, let us consider the contribution of a single decay, ${\cal R}^{(1)}$, in the simplest case, $n=0$. Since for $n=0$ only vacuum $e^{-}e^{+}$ emission is allowed, we will only consider energies $E_{d}>E_{0}^{\textrm{th}}$. If $E_{d}<E_{0}^{\textrm{th}}$ the decay width vanishes and no constraint on LIV can be derived.
@@
Higher–order contributions would involve even larger parent energies, where the source emission is strongly suppressed. In fact, from a general study of the contributions in the cascade, it is possible to show  [^2] This analysis lies beyond the scope of the present work and will be presented elsewhere. that a bound on $\mathcal{R}^{(k)}$ is just
$\mathcal{R}^{(k)}<e^{-L\Gamma(E_{d})}\left<\frac{y^{\gamma}}{1-y^{5}}\right>^{k}\,.$ (26)
Then, a bound on the quotient between the detected and emitted flux can be build as a geometric sum,
@@
V Time delays, KM3–230213A and future UHE neutrino events
The unified framework developed in Section II allows one to derive updated constraints on LIV from the detection of ultra–high–energy neutrinos. In this section we apply these results to the specific case of the KM3–230213A event recently reported by the KM3NeT collaboration, and we discuss their implications for present and future observations.
V.1 Constraints from the KM3–230213A event
The KM3–230213A event corresponds to a neutrino of energy $E_{d}\simeq 100$ PeV. Using the formalism of Sec. III and the results in Figs. 1, 2, 3 and 4, we translate this detection into constraints on superluminal LIV for both the energy–independent ($n=0$) and quadratic ($n=2$) scenarios.
  • Case $n=0$: For sources in the local Universe the bounds on $\delta$ are very close to those reported by KM3NeT: the slope of the constraint as a function of distance/redshift is practically the same, with only a modest ($\sim$20%) tightening attributable to the use of the full decay widths instead of the Cohen–Glashow approximation and the proper flavour averaging. The most relevant difference appears for cosmological sources, where the expansion is treated consistently here: the curves bend down at large redshift (cf. Fig. 1), leading to stronger limits than in Ref. [13]. In particular, for $z\in[1,3]$ we obtain
    $\delta\;\lesssim\;10^{-22}\text{--}10^{-23},$
    whereas the KM3NeT estimate stays closer to $\delta\sim 10^{-22}$ because the high-$z$ propagation is not handled consistently in their survival–probability treatment. This difference is precisely the effect of the cosmological redshifting of the energy inside the decay width.
  • Case $n=2$: For the quadratic case, the constrained parameter is the LIV scale $\Lambda$ defined by $E^{2}=p^{2}[1+(p/\Lambda)^{2}]$. From Figs. 3 and 4, and taking $E_{d}\simeq 100$ PeV, we find

### v2

# Superluminal constraints from ultra-high-energy neutrino events

@@ Abstract
The $220^{+570}_{-100}\,$PeV neutrino detected by KM3NeT marks the beginning of ultra-high-energy neutrino astronomy and provides a powerful probe of Lorentz Invariance Violation (LIV). In superluminal scenarios, neutrinos can decay through vacuum $e^{-}e^{+}$ pair emission or neutrino splitting. Previous analyses of the KM3-230213A event relied on simplified survival-probability estimates and, in some cases, used inaccurate decay-width expressions or neglected redshift and threshold effects. In this work we present a unified and self-consistent framework that corrects these issues and applies to both the energy-independent ($n=0$) and quadratic ($n=2$) superluminal cases. We collect and recast the decay-width and threshold expressions, clarify their flavor dependence, and include a consistent treatment of cosmological propagation. We also assess the impact of cascade regeneration and show that cascade effects are negligible for the purpose of setting LIV bounds. The survival-probability approximation adopted in previous works is therefore justified, while our framework provides a coherent basis for future analyses of superluminal neutrino constraints, which should consistently include possible time-delay signatures.
@@ I Introduction
The recent detection of a neutrino event with an energy of $220^{+570}_{-100}$ PeV by KM3NeT [1] has opened the window of ultra-high-energy (UHE) neutrino astronomy. While its origin—whether it arises from a point source, the diffuse background, or is a cosmogonic neutrino—remains uncertain [2, 3, 4, 5, 6, 7], its mere observation confirms the existence of a population of neutrinos surpassing the 100 PeV threshold [8]. This detection also raises intriguing questions, particularly about the absence of similar events in IceCube, which has so far only reported neutrinos up to an order of magnitude lower in energy [9]. However, it is reasonable to anticipate that this event is just the first of many within this energy regime, to be observed by IceCube, KM3NeT, and next-generation UHE neutrino observatories such as IceCube-Gen2 (radio), RNO-G, GRAND, and POEMMA, which will significantly increase the exposure to neutrinos in the EeV range.
This newly opened window provides an unprecedented opportunity to explore new physics. In fact, the particularities of the KM3-230213A event have been used to connect it with phenomena as diverse as oscillations and sterile neutrinos [10], dark matter [11, 12, 13, 14], and primordial black holes [15, 16, 17, 18, 19].
A population of UHE neutrinos have general implications for new interactions and neutrino decays. In particular, it affects scenarios of Lorentz Invariance Violation (LIV), where neutrinos with a modified dispersion relation may undergo decay during propagation, altering the expected flux of cosmic neutrinos [20].
@@
The present work revisits these analyses within a unified and self-consistent framework. We incorporate a proper treatment of the decay thresholds, the energy and flavor dependence of the decay widths, and the cosmological propagation of neutrinos. We also examine the validity and limitations of the survival-probability approach by explicitly studying the potential influence of cascade regeneration on the detected flux. This provides a consistent methodology applicable to both energy-independent and energy-dependent superluminal scenarios, and clarifies the domain of applicability of current and future constraints based on ultra-high-energy neutrino observations.
The structure of the paper is as follows. In section II we review the expressions of the decay widths relevant for different LIV scenarios, emphasizing their unified formulation and comparing them with those adopted in the literature. These results are then applied in section III to derive constraints on superluminal neutrinos using the survival-probability method, incorporating the proper redshift dependence and decay thresholds. In section IV we examine the possible impact of cascade regeneration on these results and demonstrate that its contribution is negligible for setting LIV bounds, thereby justifying the validity of the simplified approach. Section V applies this framework to the KM3-230213A event and discusses its implications for future ultra-high-energy neutrino observations, including the connection with possible time-delay effects. Finally, section VI summarizes our conclusions and outlines prospects for future studies.
@@ II Unification of decay widths for superluminal neutrinos
More accurate results for the decay width were obtained in refs. [27, 28], which showed that the original Cohen-Glashow estimate [25] follows from a prescription that cannot be derived from a consistent Lagrangian framework. More recently, ref. [26] improved the calculation of both total and partial widths by incorporating the — previously neglected — charged-current contribution to the vacuum $e^{-}e^{+}$ pair-emission (VPE) channel, and by considering that another decay process, $\nu_{\alpha}\to\nu_{\alpha}\nu_{\beta}\bar{\nu}_{\beta}$, neutrino splitting (NSpl), is allowed for $n\neq 0$. The decay widths for the processes $(i)=\{\text{VPE},\text{NSpl}\}$ can be written as
$\Gamma_{\alpha,n}^{(i)}(E)\simeq K_{\alpha,n}^{(i)}\,\frac{G_{F}^{2}}{192\pi^{3}}\,E^{5+3n}\,\Delta_{n}^{3}\,$, (6)
where the $\simeq$ symbol means that $E$ is well above the threshold of the corresponding process. The values of the constants $K_{\alpha,n}^{(i)}$ are given in tables 1 and 2, and take into account that neutrino splitting can proceed through three channels, corresponding to each possible flavor $\beta$ of the neutrino-antineutrino pair. [^1] The results in tables 1 and 2 are taken from ref. [26]. Note, however, that in order to follow the usual convention in the $n=0$ case, where $\delta$ is defined as $\delta=v^{2}-1$, the parameter $\Lambda$ in the present work is defined through the relation between $E^{2}$ and $\vec{p}^{\,2}$, whereas ref. [26] used the relation between $E$ and $|\vec{p}|$. Consequently, the high-energy scale in ref. [26] corresponds to $2^{1/n}\,\Lambda$.
@@
$\Gamma_{\alpha,n}^{\text{(All)}}(E)\simeq\Big(K_{\alpha,n}^{\text{(VPE)}}+K_{\alpha,n}^{\text{(NSpl)}}\Big)\,\frac{G_{F}^{2}}{192\pi^{3}}\,E^{5+3n}\,\Delta_{n}^{3}\,$, (8)
@@
$\frac{m_{e}^{2}/E^{2}}{E^{n}\Delta_{n}}\sim\left(\frac{E_{n}^{\mathrm{th}}}{E}\right)^{n+2}\,.$ (10)
@@
III Constraints on superluminal neutrinos from the survival-probability method
@@
Previous studies constraining superluminal neutrinos [21, 22, 23] have neglected the flavor dependence of the decay widths. In addition, they often rely on the Cohen-Glashow approximation rather than on the full results of refs. [27, 28, 26], or restrict the analysis to the case $n=0$. In particular, ref. [21] presented constraints on the LIV parameter $\delta$ for the $n=0$ case, employing the Cohen-Glashow coefficient $1/14\approx 0.071$ instead of the correct value $17/84\approx 0.202$ given in table 1. In addition, the analysis does not incorporate the VPE threshold in their figure 1, effectively assuming $E>E^{\text{th}}_{n}$, while, as they point out in the main text, the validity of this assumption is limited. Constraints are then derived on $\Delta_{0}=\delta$ as a function of the source distance $L$, using the survival probability $\mathcal{P}^{\text{SV}}_{n}$ introduced in eq. (11), which strictly applies only to local sources. Nevertheless, this approximation is extended to cosmological distances in their figure 1, up to the scale of the observable Universe. In what follows, we address the shortcomings discussed above and develop a framework that overcomes these limitations. Our aim is to provide an analysis scheme that can be applied not only to the KM3-230213A event, but also to future detections of ultra-high-energy neutrinos.
@@
Figure 1: Constraints on the LIV parameter $\delta$ as a function of the source redshift $z$ (bottom axis) and the corresponding source distance $L$ in meters (top axis). The left vertical axis shows the upper bound on $\delta$ obtained from eqs. (15) and (16) for different detected energies, each represented by a different color as indicated in the legend. The right vertical axis displays the corresponding VPE threshold energy $E^{\text{th}}_{0}$ associated with the values of $\delta$ shown on the left axis.
@@
Figure 2: Heat-map representation of the constraints in the $(E_{d},\delta,z)$ space for the $n=0$ case. The bottom axis shows the LIV parameter, while the top axis gives the corresponding VPE threshold energy, which is also shown as a dashed line in the plot. The left vertical axis corresponds to the detected energy $E_{d}$. Colors encode the constraint derived from the survival-probability condition in the range from $L=10^{17}\,\text{m}$ up to $z=10$. A detected event corresponds to a horizontal line; its intersection with the colored boundary yields the upper bound on $\delta$ at each $z$.
@@
In the $n=0$ case, the curve in figure 1 of ref. [21] is replaced by the set of upper-limit curves we show in our figure 1. These display the upper bound on $\delta$ as a function of the source redshift $z$ (bottom axis) and the corresponding comoving distance $L$ (top axis). The right vertical axis indicates the VPE threshold energy $E^{\text{th}}_{0}$ corresponding to the value of $\delta$ on the left vertical axis, and each colored curve corresponds to a different detected energy $E_{d}$. Values of $\delta$ above a given curve are excluded for a source at the corresponding $z$ and $E_{d}$. One can also see that the curves are no longer straight lines: the inclusion of cosmic expansion changes their slope at high redshift, and the resulting bound on $\delta$ becomes very strong. In fact, independently of the detected energy $E_{d}$, there always exists a sufficiently distant source ($z$ large enough) for which the survival probability amounts to $\epsilon$. The cosmological effect becomes important already at redshifts $z\sim 1$, which implies that neglecting it, as in ref. [21], leads to a physically relevant inconsistency in the resulting limits.
@@
A more powerful way of visualizing the constraints is provided by the heat-map shown in figure 2, where the three variables $(E_{d},\delta,z)$ are combined into a single representation. Its key advantage is that a detected event corresponds simply to a horizontal line at the measured energy, so that the corresponding upper bound on $\delta$ can be directly read as a function of the source redshift $z$. This representation is therefore valid not only for the present analysis but also for future UHE neutrino detections. The plot spans from a source distance $L=10^{17}\,\text{m}$ (local Universe) up to a redshift $z=10$, and the dashed line indicates the VPE threshold energy associated with each value of $\delta$.
@@
For $n=2$, the relevant parameter to be constrained is not $\delta$, but the LIV scale $\Lambda$, which we express in units of the Planck energy $E_{\text{Pl}}$. Figure 3 shows the lower bound on $\Lambda/E_{\text{Pl}}$ as a function of the source redshift $z$ (bottom axis) and the corresponding source distance $L$ (top axis), for a fixed detected energy $E_{d}$. As in the $n=0$ case, the curves deviate from straight lines at high redshift due to the effect of cosmic expansion, which substantially tightens the constraint.
Figure 3: Constraints on the LIV parameter $\Lambda$ in units of the Planck energy $E_{\text{Pl}}$ as a function of the source redshift $z$ (bottom axis) and the corresponding source distance $L$ (top axis), for fixed detected energy $E_{d}$. The left vertical axis shows the bound on $\Lambda/E_{\text{Pl}}$ obtained from the survival-probability analysis, while the right vertical axis indicates the associated VPE threshold energy. As in the $n=0$ case, the inclusion of cosmic expansion modifies the slope of the curves at large $z$, and the bound on $\Lambda$ becomes significantly stronger.
Figure 4: Heat-map representation of the constraints in the $(E_{d},\Lambda/E_{\text{Pl}},z)$ space for the $n=2$ case. The bottom axis shows the LIV parameter, while the top axis gives the corresponding VPE threshold energy, which is also shown as a dashed line in the plot. The left vertical axis corresponds to the detected energy $E_{d}$. Colors encode the constraint derived from the survival-probability condition in the range from $L=10^{17}\,\text{m}$ up to $z=10$. A detected event corresponds to a horizontal line; its intersection with the colored boundary yields the lower bound on $\Lambda$ at each $z$.
A more event-independent view is provided in figure 4, which displays a heat-map representation in the $(E_{d},\Lambda/E_{\text{Pl}},z)$ space. A detected event of energy $E_{d}$ corresponds to a horizontal line, and the intersection with the shaded region directly provides the lower bound on $\Lambda/E_{\text{Pl}}$ as a function of the source redshift. The plot spans the range $L=10^{17}\,\text{m}$ to $z=10$, and the dashed curve marks the corresponding threshold energy.
IV Validity of the survival-probability approximation
The analysis carried out so far has been implicitly based on the assumption that the detected neutrino was directly emitted by the source. In other words, the survival-probability formalism accounts only for the case in which the neutrino has propagated from the source without undergoing any decay. However, in the presence of superluminal propagation, neutrinos may decay along their trajectory, giving rise to a cascade of secondary neutrinos that can also be detected.
@@ IV Validity of the survival-probability approximation
$\phi_{d}^{(1)}(E_{d},L)=\int_{0}^{1}dy\,f(y)\,\phi_{e}\left(E_{d}/y\right)\int_{0}^{1}e^{-(1-x)L\,\Gamma(E_{d}/y)}\,L\,dx\,\Gamma(E_{d}/y)\,e^{-xL\,\Gamma(E_{d})}\,$, (18)
where $f(y)$ is the distribution of the energy fraction $y$ carried by the secondary neutrino produced in the decay. The first exponential factor accounts for the survival probability of the parent neutrino up to the decay point, the middle factor gives the decay probability within a distance interval LdxL\,dx, and the last exponential factor represents the survival of the daughter neutrino on its way to the detector.
@@
The simplified method introduced in section III, which retains only $\mathcal{R}^{(0)}$, can in principle underestimate the flux of cascade neutrinos. To assess the importance of this effect, let us consider the contribution of a single decay, ${\cal R}^{(1)}$, in the simplest case, $n=0$. Since for $n=0$ only vacuum $e^{-}e^{+}$ emission is allowed, we will only consider energies $E_{d}>E_{0}^{\textrm{th}}$. If $E_{d}<E_{0}^{\textrm{th}}$ the decay width vanishes and no constraint on LIV can be derived.
@@
Higher-order contributions would involve even larger parent energies, where the source emission is strongly suppressed. In fact, from a general study of the contributions in the cascade, it is possible to show that
$\mathcal{R}^{(k)}<e^{-L\Gamma(E_{d})}\left<\frac{y^{\gamma}}{1-y^{5}}\right>^{k}\,$, (26)
for every value of kk. [^2] This analysis lies beyond the scope of the present work and will be presented elsewhere. Then, a bound on the quotient between the detected and emitted flux can be build as a geometric sum,
@@
V Time delays, KM3-230213A and future UHE neutrino events
The unified framework developed in section II allows one to derive updated constraints on LIV from the detection of ultra-high-energy neutrinos. In this section we apply these results to the specific case of the KM3-230213A event recently reported by the KM3NeT collaboration, and we discuss their implications for present and future observations.
V.1 Constraints from the KM3-230213A event
The KM3-230213A event corresponds to a neutrino of energy $220^{+570}_{-100}$ PeV. Using the formalism of section III and the results in figures 1, 2, 3 and 4, we translate this detection into constraints on superluminal LIV for both the energy-independent ($n=0$) and quadratic ($n=2$) scenarios. To derive the most conservative constraints, we adopt the lower bound of the event’s energy interval, approximated here as $E_{d}\approx 100$ PeV.
  • Case $n=0$: For sources in the local Universe the bounds on $\delta$ are very close to those reported by KM3NeT: the slope of the constraint as a function of distance/redshift is practically the same, with only a modest ($\sim$20%) tightening attributable to the use of the full decay widths instead of the Cohen-Glashow approximation and the proper flavour averaging. The most relevant difference appears for cosmological sources, where the expansion is treated consistently here: the curves bend down at large redshift (cf. figure 1), leading to stronger limits than in ref. [21]. In particular, for $z\in[1,3]$ we obtain
    $\delta\;\lesssim\;10^{-23}\text{--}10^{-22},$
    whereas the KM3NeT estimate stays closer to $\delta\sim 10^{-22}$ because the high-$z$ propagation is not handled consistently in their survival-probability treatment. This difference is precisely the effect of the cosmological redshifting of the energy inside the decay width.
  • Case $n=2$: For the quadratic case, the constrained parameter is the LIV scale $\Lambda$ defined by $E^{2}=p^{2}[1+(p/\Lambda)^{2}]$. From figures. 3 and 4 we find